# Distributed Hierarchical Machine Learning for Joint Resource Allocation and Slice Selection in In-Network Edge Systems

### v2

@@ I Introduction
The metaverse has attracted considerable attention from academia and industry in the last few years, and significant advancements have been in this technology owing to recent developments such as extended reality, 5G, and edge intelligence, among others [10415393]. However, the extremely high demand for resources (for example, computing, networking, and storage) in the Metaverse is one of the major challenges that prevent Metaverse deployment [9944868]. An enormous number of resources is required to fulfill the Quality-of-Service (QoS) requirements and meet user experience expectations in the Metaverse. Metaverse, which can simultaneously host a hundred thousand users, is expected to support millions of users concurrently [bhattacharya2023towards].
@@
To meet these demands, mobile edge computing (MEC) has been widely adopted as a key enabler, bridging the gap between computational resources end-users to reduce latency and improve service quality [10754988]. Although MEC offers remedies through remote offloading (RO), it does not meet concurrent user demands [aliyu2023dynamic]. “Computing in the network” (COIN), sometimes termed ”in-network computing,” has emerged as a complementary approach that extends computing resources closer to the user by leveraging distributed nodes within the network [9919270]. COIN processes tasks at various points along the data path, reducing latency and improving resource utilization. The idea of COIN is not to replace the MEC, but to collaborate with the MEC [aliyu2023dynamic], leveraging additional computational resources distributed across the network. However, augmenting computing resources or enabling COIN leads to heightened power consumption [10575680]. Effectively allocating COIN resources in real-time to adapt to continually shifting user demands, while ensuring overall system availability, poses a critical challenge.
@@
Network slicing allows for the creation of virtual networks (slices) over a shared physical infrastructure, with each slice designed to meet the specific requirements of different applications [8685766]. In the context of the Metaverse, slicing can provide tailored computational, bandwidth, and latency conditions to accommodate diverse use cases (from gaming to healthcare) [10158923]. However, the dynamic allocation of resources across these slices is a challenging task; dynamic allocation ensures that resources are optimally distributed between communication and computation while adapting to the varying demands of users and tasks.
@@
Previous researchers [sasan2024joint], addressed joint network slicing and in-network computing resource allocation using a water-filling-based heuristic algorithm. However, their focus was solely on managing resources between slices (inter-slice), without considering the resource management issues within the slices (intra-slice). Furthermore, heuristic algorithms can achieve optimal solutions in a feasible amount of time, but they are not easy to derive [lia2022].
@@
Recently, AI approaches as proposed in the concept of intelligent edge and considered a key 6G enabler have been used to determine how to allocate computing and communication resources across networks [8884234]. This methodology holds promise by replacing complex mathematical modeling of the system to derive mathematically tractable heuristics with a data-driven understanding of the network. For example, the authors of [araujo2021hybrid] focused on the allocation of network resources to provide service function chains (SFCs). After proposing a linear integer programming problem, they developed a hybrid strategy combining the optimal approach with supervised machine learning (ML) to find adequate network resources while decreasing the resolution time.
@@ II-A Computing in the network (COIN) / In-Network Computing
Sapio et al[sapio2017network] describe COIN as a dumb idea whose time has come. It extends computing infrastructure closer to the user, typically within proximity to end-users or IoT devices [kunze2021investigating]. Several studies have explored the potential of COIN in various domains, including metaverse, holographic applications, IoT, smart cities, and multimedia applications [lia2022, 11050417, aliyu2023dynamic, 10575680, 10018567]. The authors in [lia2022] considered the problem of optimal placement of delay-constrained tasks to COIN nodes in an intelligent edge based on specific requirements to minimize network resource usage for low-latency task execution. Similarly, in a study conducted by [11050417], the authors developed a task-offloading solution for delay-constrained Metaverse systems to determine the optimal decision for offloading rendering tasks to COIN nodes or edge servers. The work in [10018567] proposes an SDN-based COIN framework to place reusable, delay-sensitive computing tasks directly within the data forwarding path. Aliyu et al. [10575680] propose a COIN-assisted edge architecture for Industrial IoT to reduce computation latency via partial offloading over URLLC links. The authors formulate a joint optimization of offloading decisions, ratios, and resource allocations, leveraging a hybrid game-theoretic approach combined with Double Deep Q-Networks (DDQN) to support low-complexity and distributed resource management.
@@ II-B ML-based Resource Management
Solving targeted optimization problems, such as minimizing energy consumption, latency, or bandwidth usage, either individually or jointly, typically involves addressing computationally intensive NP-hard problems. According to [8847416], machine learning (ML) techniques offer a promising alternative to traditional mathematical modeling of the system, enabling the derivation of mathematically tractable heuristics with a data-driven understanding of the system. For instance, [10041775] proposed a collaborative ML resource allocation scheme tailored for SDN-enabled fog computing, while [lia2022] analyzed four classes of supervised learning methods; Decision tree (DT), Support vector machine (SVM), Multi-layer perceptron (MLP), and Bagged trees (BT) for task placement optimization at the network edge. In [silva2024ml], the authors proposed a predictive ML-based approach to monitor and guarantee the quality of network slice service by analyzing throughput and packet loss rate key quality indicators (KQI). In [hejazi2024learning], SVM and MLP were employed for application placement in mobile edge computing servers. By formulating the problem as a two-stage stochastic optimization model, this study demonstrated the effectiveness of ML models in improving solution times and decision-making accuracy for user-to-server request allocation.
While previous works employing ML focus on instance based inference, where each decision is derived for individual users or requests in isolation, we addressed a set-structured/set input problem. We consider groups of WDs jointly, treating the task as a multiple instance learning problem. To capture this structure, we employ a DeepSet-based encoder–decoder architecture [zaheer2017deep], enabling the model to learn permutation-invariant representations of the WD sets.
@@ III System Model
In the case of assigning the task to slice, it goes through exactly one AP and can either be computed within the network on a set of COIN $\mathcal{C}=\{1,2,\ldots,C\}$, or can be offloaded to a set of MEC $\mathcal{M}=\{1,2,\ldots,M\}$. For simplicity, we denote $\overline{N}$ to be the set of all edge nodes (ENs), i.e $\mathcal{C}\cup\mathcal{M}$. The APs and ENs form the set $\mathcal{E}\triangleq\overline{N}\cup\mathcal{A}$ of edge resorces. For all the task generated, there is expected number of instructions $R_{i}$ required to perform the computation. Since the WD, COIN and MEC have different instruction set of architectures, the required number of instructions may also vary. Therefore, for a task generated by WD $i$ we denote by $\mathcal{L}_{i}$, $\mathcal{L}_{i,n}$ as the expected number of instructions locally and in slice $n$ respectively which we considered them to be estimated by methods described in [jovsilo2022joint, 8314708].
@@ III-B Computing Resources
We assumed like in the literature [11050417, lia2022], and justified by empirical studies [zukerman2013introduction], the arrival rate of requests of task generated by WD $i$ at node $j$ of slice $n$ follows a Poisson distribution with parameter $\alpha_{i}$. The expected number of instructions $\mathcal{L}_{i,n}$ required to execute a task generated by WD $i$ in slice $n$ is exponentially distributed. Therefore, all nodes can form an M/M/1 queuing model to process their corresponding computing tasks. The total request arrival rate at node $j$, which still follows a Poisson distribution, can be defined as:
@@ V Distributed Hierarchical DeepSet-S model for resource management and task offloading
The MINLP problem formulated in equation 13 corresponds to the Generalized Assignment Problem (GAP) [cattrysse1992survey], which is an NP-hard problem in the combinatorial optimization literature. Due to its complexity, resolving it by using a standard optimization solver is not feasible; it would take a long time to arrive at the optimal solution. Therefore, a more dynamic solution is needed to solve the MINLP problem efficiently, even at the expense of a suboptimal solution.
For this purpose, consistent with the recent studies [11050417, lia2022], we decompose the MINLP into a sequence of $\mathcal{N}+2$ optimization subproblems, which are solved sequentially using machine learning (ML)-based techniques.
@@ VI-B Experiment Environment
We consider a small-scale scenario for an edge application with a square area of (500 x 500)$m$ where 8 COINs, 1 MEC, 3 APs, and WDs are placed at random. The power spectral density of Gaussian noise $N_{0}$ is 174dBm/Hz-174\ dBm/Hz [aliyu2024digital], and according to [jovsilo2022joint], we represent the achievable physical rate $R_{i,a}$ from WD $i$ to AP $a$ as $R_{i,a}=B_{a}log\ (1+(distance)P^{t}/N_{0})$, where distancedistance is the Euclidean distance between WD $i$ and AP $a$, and $P^{t}$ is the transmission power of WD $i$. Other simulation parameters are summarized in table II.
@@
COINs node capacity $[5,10]$ x $10^{8}$ CPUcycles/sCPU\ cycles/s [lia2022]
MEC capacity 11 x $10^{10}$ CPUcycles/sCPU\ cycles/s [lia2022]
WD local computing capacity 22 x $10^{7}$ CPUcycles/sCPU\ cycles/s [sardellitti2015joint]
Input task size $[1,10]$ MB [aliyu2024digital]
Bandwidth $B_{a}$ per AP 18 MHz [ahmadi2009overview]
Transmit Power of WD’s $10^{-6},0.1]W [$jovsilo2022joint]
@@
In Table III, we evaluate the performance of our DeepSets-S model against its variant DeepSets which disables slack, and apply normalization on the data to satisfy the resource allocation constraint, and other state-of-the-art multi-instance learning model; SetTransformer [lee2019set]. We evaluate using Acc@k, where Acc@0.5, Acc@1, and Acc@2 measure the proportion of predictions that fall within 0.5, 1, and 2 units of the ground truth, respectively.
@@
  • Optimization Solvers (OS): We solve SP1 and SP2 through Ipopt optimization solver [wachter2006implementation], and SP3 through Gurobi optimization solver [gurobi] to achieve the optimal offloading decision vector $\boldsymbol{\delta}$.

### v3

@@ I Introduction
The metaverse has attracted considerable attention from academia and industry in the last few years, and significant advancements have been in this technology owing to recent developments such as extended reality, 5G, and edge intelligence, among others [35]. However, the extremely high demand for resources (for example, computing, networking, and storage) in the Metaverse is one of the major challenges that prevent Metaverse deployment [33]. An enormous number of resources is required to fulfill the Quality-of-Service (QoS) requirements and meet user experience expectations in the Metaverse. Metaverse, which can simultaneously host a hundred thousand users, is expected to support millions of users concurrently [7].
@@
To meet these demands, mobile edge computing (MEC) has been widely adopted as a key enabler, bridging the gap between computational resources end-users to reduce latency and improve service quality [14]. Although MEC offers remedies through remote offloading (RO), it does not meet concurrent user demands [4]. “Computing in the network” (COIN), sometimes termed ”in-network computing,” has emerged as a complementary approach that extends computing resources closer to the user by leveraging distributed nodes within the network [16]. COIN processes tasks at various points along the data path, reducing latency and improving resource utilization. The idea of COIN is not to replace the MEC, but to collaborate with the MEC [4], leveraging additional computational resources distributed across the network. However, augmenting computing resources or enabling COIN leads to heightened power consumption [3]. Segal et al. [28] have shown the potential of in-network computing for reducing network utilization cost; however, effectively allocating COIN resources in real-time to adapt to continually shifting user demands, while ensuring overall system availability, poses a critical challenge.
@@
Network slicing allows for the creation of virtual networks (slices) over a shared physical infrastructure, with each slice designed to meet the specific requirements of different applications [36, 9]. In the context of the Metaverse, slicing can provide tailored computational, bandwidth, and latency conditions to accommodate diverse use cases (from gaming to healthcare) [10]. However, the dynamic allocation of resources across these slices is a challenging task; dynamic allocation ensures that resources are optimally distributed between communication and computation while adapting to the varying demands of users and tasks.
@@
Previous researchers [27], addressed joint network slicing and in-network computing resource allocation using a water-filling-based heuristic algorithm. However, their focus was solely on managing resources between slices (inter-slice), without considering the resource management issues within the slices (intra-slice). Furthermore, heuristic algorithms can achieve optimal solutions in a feasible amount of time, but they are not easy to derive [19].
@@
Recently, AI approaches as proposed in the concept of intelligent edge and considered a key 6G enabler have been used to determine how to allocate computing and communication resources across networks [11]. This methodology holds promise by replacing complex mathematical modeling of the system to derive mathematically tractable heuristics with a data-driven understanding of the network. For example, the authors of [6] focused on the allocation of network resources to provide service function chains (SFCs). After proposing a linear integer programming problem, they developed a hybrid strategy combining the optimal approach with supervised machine learning (ML) to find adequate network resources while decreasing the resolution time.
@@ II-A Computing in the network (COIN) / In-Network Computing
Sapio et al[25] describe COIN as a dumb idea whose time has come. It extends computing infrastructure closer to the user, typically within proximity to end-users or IoT devices [17]. Several studies have explored the potential of COIN in various domains, including metaverse, holographic applications, IoT, smart cities, and multimedia applications [19, 21, 4, 3, 5, 32]. The authors in [19] considered the problem of optimal placement of delay-constrained tasks to COIN nodes in an intelligent edge based on specific requirements to minimize network resource usage for low-latency task execution. Similarly, in a study conducted by [21], the authors developed a task-offloading solution for delay-constrained Metaverse systems to determine the optimal decision for offloading rendering tasks to COIN nodes or edge servers. The work in [5] proposes an SDN-based COIN framework to place reusable, delay-sensitive computing tasks directly within the data forwarding path. Aliyu et al. [3] propose a COIN-assisted edge architecture for Industrial IoT to reduce computation latency via partial offloading over URLLC links. The authors formulate a joint optimization of offloading decisions, ratios, and resource allocations, leveraging a hybrid game-theoretic approach combined with Double Deep Q-Networks (DDQN) to support low-complexity and distributed resource management. Wu et al. [32] optimized COIN performance by adjusting the Computing Data Unit (CDU) size to reduce end-to-end latency in multi-hop in-network computing systems.
@@ II-B ML-based Resource Management
Solving targeted optimization problems, such as minimizing energy consumption, latency, or bandwidth usage, either individually or jointly, typically involves addressing computationally intensive NP-hard problems. According to [24], machine learning (ML) techniques offer a promising alternative to traditional mathematical modeling of the system, enabling the derivation of mathematically tractable heuristics with a data-driven understanding of the system. For instance, [30] proposed a collaborative ML resource allocation scheme tailored for SDN-enabled fog computing, while [19] analyzed four classes of supervised learning methods; Decision tree (DT), Support vector machine (SVM), Multi-layer perceptron (MLP), and Bagged trees (BT) for task placement optimization at the network edge. In [29], the authors proposed a predictive ML-based approach to monitor and guarantee the quality of network slice service by analyzing throughput and packet loss rate key quality indicators (KQI). Liu et al. [20] proposed a learning-assisted end-to-end network slicing framework for resource allocation and task completion time minimization in heterogeneous networks. In [13], SVM and MLP were employed for application placement in mobile edge computing servers. By formulating the problem as a two-stage stochastic optimization model, this study demonstrated the effectiveness of ML models in improving solution times and decision-making accuracy for user-to-server request allocation.
While previous works employing ML focus on instance based inference, where each decision is derived for individual users or requests in isolation, we addressed a set-structured/set input problem. We consider groups of WDs jointly, treating the task as a multiple instance learning problem. To capture this structure, we employ a DeepSet-based encoder–decoder architecture [34], enabling the model to learn permutation-invariant representations of the WD sets.
@@ III System Model
In the case of assigning the task to slice, it goes through exactly one AP and can either be computed within the network on a set of COIN $\mathcal{C}=\{1,2,\ldots,C\}$, or can be offloaded to a set of MEC $\mathcal{M}=\{1,2,\ldots,M\}$. For simplicity, we denote $\overline{N}$ to be the set of all edge nodes (ENs), i.e $\mathcal{C}\cup\mathcal{M}$. The APs and ENs form the set $\mathcal{E}\triangleq\overline{N}\cup\mathcal{A}$ of edge resorces. For all the task generated, there is expected number of instructions $R_{i}$ required to perform the computation. Since the WD, COIN and MEC have different instruction set of architectures, the required number of instructions may also vary. Therefore, for a task generated by WD $i$ we denote by $\mathcal{L}_{i}$, $\mathcal{L}_{i,n}$ as the expected number of instructions locally and in slice $n$ respectively which we considered them to be estimated by methods described in [15, 22].
@@ III-B Computing Resources
We assumed like in the literature [21, 19], and justified by empirical studies [37], the arrival rate of requests of task generated by WD $i$ at node $j$ of slice $n$ follows a Poisson distribution with parameter $\alpha_{i}$. The expected number of instructions $\mathcal{L}_{i,n}$ required to execute a task generated by WD $i$ in slice $n$ is exponentially distributed. Therefore, all nodes can form an M/M/1 queuing model to process their corresponding computing tasks. The total request arrival rate at node $j$, which still follows a Poisson distribution, can be defined as:
@@ V Distributed Hierarchical DeepSet-S model for resource management and task offloading
The MINLP problem formulated in equation 13 corresponds to the Generalized Assignment Problem (GAP) [8], which is an NP-hard problem in the combinatorial optimization literature. Due to its complexity, resolving it by using a standard optimization solver is not feasible; it would take a long time to arrive at the optimal solution. Therefore, a more dynamic solution is needed to solve the MINLP problem efficiently, even at the expense of a suboptimal solution.
For this purpose, consistent with the recent studies [21, 19], we decompose the MINLP into a sequence of $\mathcal{N}+2$ optimization subproblems, which are solved sequentially using machine learning (ML)-based techniques.
@@ VI-B Experiment Environment
We consider a small-scale scenario for an edge application with a square area of (500 x 500)$m$ where 8 COINs, 1 MEC, 3 APs, and WDs are placed at random. The power spectral density of Gaussian noise $N_{0}$ is 174dBm/Hz-174\ dBm/Hz [2], and according to [15], we represent the achievable physical rate $R_{i,a}$ from WD $i$ to AP $a$ as $R_{i,a}=B_{a}log\ (1+(distance)P^{t}/N_{0})$, where distancedistance is the Euclidean distance between WD $i$ and AP $a$, and $P^{t}$ is the transmission power of WD $i$. Other simulation parameters are summarized in table II.
@@
COINs node capacity $[5,10]$ x $10^{8}$ CPUcycles/sCPU\ cycles/s [19]
MEC capacity 11 x $10^{10}$ CPUcycles/sCPU\ cycles/s [19]
WD local computing capacity 22 x $10^{7}$ CPUcycles/sCPU\ cycles/s [26]
Input task size $[1,10]$ MB [2]
Bandwidth $B_{a}$ per AP 18 MHz [1]
Transmit Power of WD’s $10^{-6},0.1]W [$15]
@@
In Table III, we evaluate the performance of our DeepSets-S model against its variant DeepSets which disables slack, and apply normalization on the data to satisfy the resource allocation constraint, and other state-of-the-art multi-instance learning model; SetTransformer [18]. We evaluate using Acc@k, where Acc@0.5, Acc@1, and Acc@2 measure the proportion of predictions that fall within 0.5, 1, and 2 units of the ground truth, respectively.
@@
  • Optimization Solvers (OS): We solve SP1 and SP2 through Ipopt optimization solver [31], and SP3 through Gurobi optimization solver [12] to achieve the optimal offloading decision vector $\boldsymbol{\delta}$.